OpenStudy (anonymous):

Divide (18x^3+12x-3x)/6x^2

OpenStudy (lgbasallote):

what is 18x^3 divided by 6x^2?

OpenStudy (anonymous):

3x?

OpenStudy (lgbasallote):

good! can you divide 12x + 3 by 6x^2?

OpenStudy (anonymous):

2x+1/2x

OpenStudy (anonymous):

ur very close hint ^3 thats missing and at the beginning it shd be 12

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

12x+ 1/2x^3 is what i got

OpenStudy (anonymous):

oh. okay well for my anser i got 3x+2-1/2x

OpenStudy (anonymous):

alright =)

OpenStudy (anonymous):

but idk if its right

OpenStudy (lgbasallote):

Hint: if you expand this \(\LARGE \frac{18x^3}{6x^2} + \frac{12x}{6x^2} + \frac{3}{6x}\) \(\LARGE 3x + \frac{2}{x} + \frac{1}{2x}\) you were close though :D

OpenStudy (anonymous):

Okay thanks :)

OpenStudy (lgbasallote):

you know how that was got right?

OpenStudy (anonymous):

Yeah I see it.

OpenStudy (lgbasallote):

goodie